OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). sin 2x - sin 4x = 0

OpenStudy (freckles):

\[\sin(4x) \neq \sin(2x)+\sin(2x)\]

OpenStudy (chillout):

You need to use \(\sin(a+b)=\sin(a)\cos(b)+\sin(b)\cos(a)\)

OpenStudy (freckles):

\[\sin(2x)-\sin(4x) =0 \\ \text{ use double angle identity on the } \sin(4x) \\ \sin(2 \cdot [2x])=2 \sin(2x)\cos(2x) \\ \sin(2x)-2 \sin(2x) \cos(2x) =0 \\ \text{ factor out } \sin(2x) \\ \sin(2x)[1-2\cos(2x)]=0 \\ \text{ set both factors equal to 0} \\ \sin(2x)=0 \text{ or } 1-2\cos(2x)=0\]

OpenStudy (freckles):

solve both equations

OpenStudy (anonymous):

wait, I'm so confused. here are the options i have A) π/6, π/2, 5π/6, 7π/6, 3π/2, 11π/6 B) 0, π/6, π/2, 5π/6, π, 7π/6, 3π/2, 11π/6 C) 0, 2π/3, 4π/3 D) 0, π/3, 2π/3, π, 4π/3, 5π/3

OpenStudy (freckles):

on which part are you confused?

OpenStudy (anonymous):

so the answer is zero?

OpenStudy (freckles):

that is only one solution

OpenStudy (freckles):

there are others

OpenStudy (freckles):

can I ask you what you are confused on/

OpenStudy (freckles):

my work or solving the equations I asked you to solve?

OpenStudy (anonymous):

ok, so I started by making an equation 2x-4x=0

OpenStudy (freckles):

where does that equation come from?

OpenStudy (anonymous):

i made sinx=x

OpenStudy (freckles):

is there anyway you can tell me what you are confused on? was it my work or me asking you to solve those two equations I asked you to solve?

OpenStudy (freckles):

sin(x) is only approximately x for values of x near 0 you cannot use this approximation to solve your equation

OpenStudy (anonymous):

i was confused by your work

OpenStudy (freckles):

ok do you know the double angle identities? like for example what does sin(2u)=?

OpenStudy (anonymous):

2sin u cos u

OpenStudy (freckles):

sin(2u)=2sin(u)cos(u) so if u=2x then we have \[\sin(2[2x])=2\sin(2x)\cos(2x)\]

OpenStudy (freckles):

\[2 \cdot 2 =4 \\ \text{ so } \sin(4x)=\sin(2[2x])=2\sin(2x)\cos(2x) \\ \text{ so I replaced } \sin(4x) \text{ with } 2\sin(2x)\cos(2x)\]

OpenStudy (anonymous):

oh ok so the answer would be B?

OpenStudy (freckles):

\[\sin(2x)-\sin(4x)=0 \\ \sin(2x)-2 \sin(2x) \cos(2x)=0 \\ \text{ next I noticed a common factor in both terms on the left } \\ \color{red}{\sin(2x)}-2 \color{red}{\sin(2x)}\cos(2x)=0\] if you want to put a 1 next to sin(2x) you can you can do this since 1*sin(2x) is sin(2x) \[1 \cdot \color{red}{\sin(2x)}-2 \color{red}{\sin(2x)}\cos(2x)=0\] now factor out the thing in red \[\color{red}{\sin(2x)}[1-2 \cos(2x)]=0 \\ \text{ so we have to solve both of the following equations } \\ \sin(2x)=0 \text{ or } 1-2\cos(2x)=0\] if you want to replace 2x with u and solve for u first then you can come back later and solve for x but replacing u back with 2x \[\sin(u)=0 \text{ or } 1-2 \cos(u)=0 \\ \sin(u) =0 \text{ or } \cos(u)=\frac{1}{2}\] on the unit circle when do you have the y-coordinates are 0? on the unit circle when do you have the x-coordinates are 1/2?